OpenStudy (anonymous):

solve the proportion x/ x+1 = 4/ x+4

OpenStudy (sw050399):

Well, you could cross-mulitply.

OpenStudy (sw050399):

You know, x times x + 4 equals 4 times x + 1.

OpenStudy (sw050399):

Visually it looks like: x(x + 4) = 4(x + 1)

OpenStudy (anonymous):

x(x+4) = 4(x+1) x^2 + 4x = 4x + 1 x^2 - 1 = 0 (x-1)(x+1) x=1 or x=-1